Axel leaves his house, drives 23 km due North, then turns onto a street and continues in a direction 60° N of E for 40 km and finally turns onto the highway due East for 20 km. What is his total displacement from his house?



1
Expert's answer
2021-09-28T17:12:40-0400

The movement of Axel can be traced as shown in the figures below;










Taking O as the initial point, first displacement, "d_1=(23km)y" (O to A)

Second displacement, "d_2 =(40cos60x+40sin60y)" (A to B)

Third displacement, "d_3=(20km)x" (B to C)

Total displacement, "d=d_1+d_2+d_3=23y+40cos60x+40sin60y+20x"

"=(20+40cos60)x+(23+40sin60)y=(40x+57.6y)km"





Magnitude, "d=\sqrt{40^2+57.6^2}=70Km", "angle \space \theta =tan^{-1}(57.6/40)=55^o"

So d=70km, "55^o" North of East